If a square of side 8 is inscribed in a circle, what is the area of the circle?
A. 64 π B. 32 π C. 128 π D. 64 E. 32

Answer by richwmiller(17219)

the diagonal of the square will be the diameter
2*8^2=d^2
128=d^2
d=sqrt(128)8 sqrt(2)
r=4 sqrt(2)
a=pi*r^2
a=pi*(4sqrt(2))^2
a=pi*16*2=32pi
B. 32pi